OpenStudy (anonymous):

At the grocery store, you buy a carton of milk for $1.95, a loaf of bread for $0.85, and a box of cereal for $2.05. Use mental math to find the total cost of the groceries

OpenStudy (anonymous):

$5.85 1.95+2.05 = 5, 5+0.85 = 5.85

OpenStudy (anonymous):

1.95+0.85+2.05=$4.85

OpenStudy (anonymous):

Correct. This why I usually rely on Mathematica to avoid mental calculations.